Rewrite each equation in slope-intercept form. Then graph it.
2/3x-1/2y=2

Answer by rfer(12731)

2/3x-1/2y=2
-1/2y=-2/3x+2
y=(-2/3x+2)/-1/2
y=(-2/3*-2/1)+(2/1*-2/1)
y=4/3x-4